OpenStudy (anonymous):

Exponential function : y=ae^-bx for 0 ≤ x ≤ 10 Two points given : (0,6) (10, y) a) Determine the value of a Is it 6???? b) If y-60 when x=4, find the value of b using your value of 'a' from a) c) Find the height of y from the point (10,y)

jhonyy9 (jhonyy9):

so you know that e^-b = 1/e^b then you can rewriting this like y= ax/(e^b) so from (0,6) you know that y=6 when x=0 i think that i have understood it wrong the bx wan being exponent of e ,right ? so then y= a/(e^(bx) a). y*e^(bx)=a so result 6*e^(b*0) = a so 6*1=a then result a=6 b).y=60 when x=4 and a=6 so b=? 60= 6/e^(b*4) 60*e^(b*4) = 6 10*e^(b*4)= 1 do you can continue it ? i need to go hope this will help you little bye